OpenStudy (anonymous):

f(x)=-(x+1)^2+4 find the zeroes

OpenStudy (anonymous):

\[-(x+1)^2+4=0\\ (x+1)^2=4\] would be a good first step

OpenStudy (anonymous):

We then do the square root?

OpenStudy (anonymous):

yes, don't forget the \(-\) as well

OpenStudy (anonymous):

Wait, but isn't 0-4=-4

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

I know one of the zeroes are going to be one, but what about the other?

OpenStudy (anonymous):

hold on lets back up

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

\[-(x+1)^2+4=0\] add \((x+1)^2\) to both sides, that is what gives \[4=(x+1)^2\]

OpenStudy (anonymous):

i just wrote it as \[(x+1)^2=4\] if you subtract \(4\) from both sides you get \[-(x+1)^2=-4\] like you said, but you then change the signs and get \[(x+1)^2=4\]

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

that means \[x+1=2\] which gives \(x=1\) or \[x+1=-2\] which gives \(x=-3\)

OpenStudy (anonymous):

I see